OpenStudy (anonymous):

A field of grass lasts100 cows 20 days . About how many days will it last for 250 cows?

OpenStudy (earthcitizen):

50

OpenStudy (earthcitizen):

50 days

OpenStudy (anonymous):

nope

OpenStudy (earthcitizen):

75 days

OpenStudy (earthcitizen):

35 days

OpenStudy (turingtest):

keep guessing...

OpenStudy (earthcitizen):

alryt..

OpenStudy (earthcitizen):

7 days

OpenStudy (turingtest):

Still no, but at least your moving in the right direction. It should make sense that more cows means less time is needed to eat the grass...

OpenStudy (earthcitizen):

5 days

OpenStudy (turingtest):

and how did you come to that conclusion?

OpenStudy (earthcitizen):

100/20=250/x, x=50

OpenStudy (earthcitizen):

250/50 =3

OpenStudy (turingtest):

formulation is always better than guessing :)

OpenStudy (turingtest):

of course I don't know what you did above... 20/100=x/250 x=5

OpenStudy (earthcitizen):

250/50 = 5

OpenStudy (earthcitizen):

is it 5 days ?

OpenStudy (turingtest):

yes, 5 days You set up the problem correctly.

OpenStudy (turingtest):

20/100=x/250

OpenStudy (earthcitizen):

in the previous thread was it really 17.5 units approx.

OpenStudy (turingtest):

That was a rather different question...

OpenStudy (earthcitizen):

yh, i know

OpenStudy (earthcitizen):

i think i'm drunk

OpenStudy (earthcitizen):

intoxicated rather

OpenStudy (earthcitizen):

sorry, i'm sleepy nw. It's been a gr8! xmas! xx happy Boxing Day! xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx

OpenStudy (turingtest):

cheers

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

it's not 5

OpenStudy (turingtest):

No, it's 8 :/ 100 cows eat the same in 20 days that 250 do in x days, so 100(20)=250x x=8 that's what I get for staying up too late.

OpenStudy (anonymous):

thanx